OpenStudy (anonymous):

what # an be decreased by 6 and the result then doubled to produce 24?

OpenStudy (mathstudent55):

Let number = x decreased by 6 = x - 6 result doubled = 2(x - 6) 2(x - 6) = 24 Solve for x

OpenStudy (anonymous):